# Breaking the Low-Rank Dilemma of Linear Attention

### v1

@@ Abstract
The Softmax attention mechanism in Transformer models is notoriously computationally expensive, particularly due to its quadratic complexity, posing significant challenges in vision applications. In contrast, linear attention provides a far more efficient solution by reducing the complexity to linear levels. However, compared to Softmax attention, linear attention often experiences significant performance degradation. Our experiments indicate that this performance drop is due to the low-rank nature of linear attention’s feature map, which hinders its ability to adequately model complex spatial information. In this paper, to break the low-rank dilemma of linear attention, we conduct rank analysis from two perspectives: the KV buffer and the output features. Consequently, we introduce Rank-Augmented Linear Attention (RALA), which rivals the performance of Softmax attention while maintaining linear complexity and high efficiency. Based on RALA, we construct the Rank-Augmented Vision Linear Transformer (RAVLT). Extensive experiments demonstrate that RAVLT achieves excellent performance across various vision tasks. Specifically, without using any additional labels, data, or supervision during training, RAVLT achieves an 84.4% Top-1 accuracy on ImageNet-1k with only 26M parameters and 4.6G FLOPs. This result significantly surpasses previous linear attention mechanisms, fully illustrating the potential of RALA. Code will be available at https://github.com/qhfan/RALA.
@@ 1 Introduction
However, applying Transformers to vision tasks is not straightforward. Self-attention, as the core mechanism of Transformers, has quadratic complexity. As the number of tokens increases, the computational load of self-attention grows significantly, making Transformers challenging to apply to high-resolution images. Many works attempt to address this issue, with some approaches grouping vision tokens to limit the number of tokens each token can interact with [43, 11, 70, 42, 16, 58, 54]. While this method effectively reduces the model’s complexity, it sacrifices the Transformer’s crucial ability to perceive information globally. Another line of work attempts to downsample the keys and values in Softmax attention to capture global features with lower complexity [56, 57, 15, 26]. However, this approach sacrifices the model’s fine-grained perception capabilities.
@@
Despite having a form and advantages very similar to Softmax attention, the actual performance of linear attention is less than satisfactory. Although it is highly efficient, its performance shows a significant gap compared to Softmax attention. To identify the cause of this performance gap, we conduct a comparison based on the DeiT-T architecture, merely replacing different attention mechanisms (vanilla linear attention, Enhanced Linear Attention [2], Efficient Attention [50], Focused Linear Attention [23]), with results shown in Fig. 2. We perform rank analysis on the features of single heads from different feature maps ($N=196$, $d=64$, with the full rank of the matrix being 64). We find that the rank of output features from existing linear attentions are significantly lower than those from Softmax attention, indicating that the diversity of features they learn is comparatively poorer.
@@ 2 Related works
Linear attention uses kernel functions to approximate the Softmax function, allowing it to change the computation order of $Q$, $K$, and $V$ in vanilla Softmax attention, thereby reducing the computational complexity from quadratic to linear. Although the efficiency of linear attention is significantly improved compared to vanilla Softmax attention, its performance still lags considerably behind. Many methods have been proposed to improve this [45, 23, 24, 63, 68, 50]. Specifically, FLatten-Transformer [23] proposes focused function to improve the focus ability of linear attention. MLLA [24] connects the linear attention with Mamba [19]. Most of other methods start by using kernel functions to approximate Softmax or introduce DWConv to increase feature diversity. However, all above methods do not address the inherent low-rank issue of linear attention, which is shown in the Fig. 2 in Introduction.
@@ 3.2 Rank-Augmented Linear Attention
To increase the rank of the KV buffer, we introduce operations into its calculation to eliminate linear dependencies between different rows and columns. Beyond the feature diversity, the rank of the KV buffer is also closely related to the amount of information it contains, increasing the information content in the KV buffer will simultaneously raise its rank. Based on this consideration, rather than assigning a weight of 1 to each $\kappa(K_{j})^{T}V_{j},j\in[1,N]$ as in other linear attentions [50, 23], we aim to highlight the significance of each token in the KV buffer by assigning higher weights to the tokens contain more information. That is:
@@
To determine the value of $\alpha_{j}$, we draw inspiration from the design of Softmax attention. In Softmax attention, the attention scores between tokens are generally considered to represent the degree of correlation between them. Tokens more correlated with the global token are considered to contain more important information[39]. Based on fact, by calculating the attention score of the global query to each key, we determine the proportion of each token in the KV buffer, resulting in the following:
@@
As shown in Fig. 4, after introducing $\alpha_{j}$ as the modulation coefficient for $\kappa(K_{j})^{T}V_{j}$, the information content in the KV buffer increases, which significantly enhances its rank.
@@
Figure 6: Illustration of the RAVLT’s block (left) and RALA (right)
@@
where $\odot$ denotes the Hadamard product. As shown in the Fig. 5, After introducing the modulation coefficients, the feature diversity of the output feature matrix is significantly enhanced, transforming it from a low-rank state to a full-rank state. This indicates that token-specific modulation adds a degree of feature diversity to the output features.
@@ 4 Experiments
We conduct extensive experiments on multiple vision tasks, such image classification, object detection, instance segmentation, and semantic segmentation. We also make ablation studies to validate the importance of each components in RALA. More details can be found in Appendix.
@@ 4.1 Image Classification
Table 3: Comparison with the state-of-the-art on ImageNet-1K classification
@@ Results.
4.2 Inference Speed
4.3 Object Detection and Instance Segmentation
Settings.
Results.
@@ 4.5 Comparison with Other Linear Attention
Table 8: Comparison of different linear attention based on DeiT-T, Swin-T, and Swin-S. RALA surpasses its counterparts by a large margin.
@@ 4.6 Ablation Study
Table 9: Ablation on each module based on RAVLT-T. The results show that performance improvements stem from our RALA’s augmentation of matrix’s rank, rather than the selection of specific functions.
Figure 8: Analysis of the rank across different layers of the model. The result shows that RALA’s augmentation of the KV buffer and output feature effectively increases the rank of the output features at each layer.
@@ 5 Conclusion
In this paper, we analyze the reasons for the performance degradation of linear attention and conclude that the low rank of its output features is the primary factor. To address this, we propose Rank-Augmented Linear Attention (RALA). This method augments two critical computation steps in linear attention: the calculation of the KV buffer and the computation of the output features, by employing two transformations. These augmentations effectively increase the rank of the matrices, thus mitigating the low-rank issue in linear attention. Based on RALA, we construct Rank-Augmented Vision Linear Transformer (RAVLT), a novel architecture based on linear attention. Its performance surpasses many Transformers, demonstrating significant potential for application.

### v3

@@ Abstract
The Softmax attention mechanism in Transformer models is notoriously computationally expensive due to its quadratic complexity, posing significant challenges in vision applications. In contrast, linear attention offers a far more efficient solution by reducing the complexity to linear levels. However, linear attention often suffers significant performance degradation compared to Softmax attention. Our experiments indicate that this performance drop stems from the low-rank nature of linear attention’s output feature map, which hinders its ability to adequately model complex spatial information. To address this low-rank dilemma, we conduct rank analysis from two perspectives: the KV buffer and the output features. Consequently, we introduce Rank-Augmented Linear Attention (RALA), which rivals the performance of Softmax attention while maintaining linear complexity and high efficiency. Building upon RALA, we construct the Rank-Augmented Vision Linear Transformer (RAVLT). Extensive experiments demonstrate that RAVLT achieves excellent performance across various vision tasks. Specifically, without using any additional labels, data, or supervision during training, RAVLT achieves an 84.4% Top-1 accuracy on ImageNet-1k with only 26M parameters and 4.6G FLOPs. This result significantly surpasses previous linear attention mechanisms, fully illustrating the potential of RALA. Code will be available at https://github.com/qhfan/RALA.
@@ 1 Introduction
However, applying Transformers to vision tasks is not straightforward. Self-attention, as the core mechanism of Transformers, has quadratic complexity. As the number of tokens increases, the computational load of self-attention grows significantly, making Transformers challenging to apply to vision tasks. Many works attempt to address this issue, with some approaches grouping vision tokens to limit the number of tokens each token can interact with [45, 12, 74, 44, 17, 60, 56]. While this method effectively reduces the model’s complexity, it sacrifices the Transformer’s crucial ability to perceive information globally. Another line of work attempts to downsample the keys and values in Softmax attention to capture global features with lower complexity [58, 59, 16, 28]. However, this approach sacrifices the model’s fine-grained perception capabilities.
@@
Despite having a form and advantages very similar to Softmax attention, the actual performance of linear attention is less than satisfactory. Although it is highly efficient, its performance shows a significant gap compared to Softmax attention. To identify the cause of this performance gap, we conduct a comparison based on the DeiT-T architecture, merely replacing different attention mechanisms (vanilla linear attention, Enhanced Linear Attention [2], Efficient Attention [52], Focused Linear Attention [25]), with results shown in Fig. 2. We perform rank analysis on the features of single heads from different output feature maps ($N=196$, $d=64$, with the full rank of the matrix being 64). We find that the rank of output features from existing linear attentions are significantly lower than those from Softmax attention, indicating that the diversity of features they learn is comparatively poorer.
@@ 2 Related works
Linear attention uses kernel functions to approximate the Softmax function, allowing it to change the computation order of $Q$, $K$, and $V$ in vanilla Softmax attention, thereby reducing the computational complexity from quadratic to linear. Although the efficiency of linear attention is significantly improved compared to vanilla Softmax attention, its performance still lags considerably behind. Many methods have been proposed to improve this [47, 25, 26, 65, 70, 52]. Specifically, FLatten-Transformer [25] proposes focused function to improve the focus ability of linear attention. MLLA [26] connects the linear attention with Mamba [21]. Most of other methods start by using kernel functions to approximate Softmax or introduce DWConv to increase feature diversity. However, the low-rank issue still exists, as shown in Fig. 2.
@@ 3.2 Rank-Augmented Linear Attention
To increase the rank of the KV buffer, we introduce operations into its calculation to eliminate linear dependencies between different rows and columns. Beyond the feature diversity, the rank of the KV buffer is also closely related to the amount of information it contains, increasing the information content in the KV buffer will simultaneously raise its rank. Based on this consideration, rather than assigning a weight of 1 to each $\kappa(K_{j})^{T}V_{j},j\in[1,N]$ as in other linear attentions [52, 25], we aim to highlight the significance of each token in the KV buffer by assigning higher weights to the tokens that contain more information. That is:
@@
To determine the value of $\alpha_{j}$, we draw inspiration from the design of Softmax attention. In Softmax attention, the attention scores between tokens are generally considered to represent the degree of correlation between them. Tokens more correlated with the global token are considered to contain more important information[41]. Based on this fact, by calculating the attention score of the global query to each key, we determine the proportion of each token in the KV buffer, resulting in the following:
@@
As shown in Fig. 4, after introducing $\alpha_{j}$ as the modulation coefficient for $\kappa(K_{j})^{T}V_{j}$, the information content in the KV buffer increases, which significantly enhances its rank. More analysis about why the introduction of $\alpha_{j}$ can increase the matrix’s rank can be found in Appendix.
@@
Figure 6: RAVLT’s block (left) and RALA (right). We employ serial numbers to represent the sequence of calculations. $\odot$ represents Hadamard product, and $\times$ represents matrix multiplication.
@@
where $\odot$ denotes the Hadamard product. As shown in the Fig. 5, After introducing the modulation coefficients, the feature diversity of the output feature matrix is significantly enhanced, transforming it from a low-rank state to a full-rank state. This indicates that token-specific modulation adds a degree of feature diversity to the output features. More analysis can be found in the Appendix.
@@ 4 Experiments
We conduct extensive experiments on multiple vision tasks, such as image classification, object detection, instance segmentation, and semantic segmentation. We also make ablation studies to validate the importance of each components in RALA. More details can be found in Appendix.
@@ 4.1 Image Classification
Table 3: Comparison with the state-of-the-art on ImageNet-1K classification. We use ”CNN” to refer to convolutional neural networks, ”Trans” to refer to Vision Transformers, and ”Linear” to refer to models based on linear operators.
Results.
4.2 Inference Speed
4.3 Object Detection and Instance Segmentation
Settings.
@@ 4.5 Comparison with Other Linear Attention
Table 8: Comparison of different linear attention based on DeiT-T, Swin-T, and Swin-S. RALA surpasses others by a large margin.
@@ 4.6 Ablation Study
Table 9: Ablation on RAVLT-T. The results show that performance improvements stem from our RALA’s augmentation of matrix’s rank, rather than the selection of specific functions.
@@
Figure 8: Analysis of the rank across different model layers. RALA’s augmentation of the KV buffer and output feature effectively increases the rank of the output features at each layer.
@@ 5 Conclusion
In this paper, we investigate the causes of performance degradation in linear attention and conclude that the low rank of its output features is the primary factor. To address this issue, we propose Rank-Augmented Linear Attention (RALA). This method enhances two critical computation steps in linear attention—the calculation of the KV buffer and the computation of the output features—by employing two transformations. These augmentations effectively increase the rank of the matrices, thereby mitigating the low-rank problem in linear attention. Building upon RALA, we develop the Rank-Augmented Vision Linear Transformer (RAVLT), a novel architecture based on linear attention. RAVLT outperforms many existing Transformers, demonstrating significant potential for practical applications.
Appendix A More Analysis About $\alpha$
In this section, we further analyze the rank-boosting effect of $\alpha$ on the KV buffer and conduct additional ablation experiments on $\alpha$ to validate the rationale behind our design.
An analysis of the rank-boosting effect.
Assume in $\kappa(K_{j}^{T})V_{j},j\in[0,N]$ that the two components $\kappa(K_{1}^{T})V_{1}$ and $\kappa(K_{2}^{T})V_{2}$ are linearly correlated. That means:
$\kappa(K_{1}^{T})V_{1}=\lambda\kappa(K_{2}^{T})V_{2}.$ (7)
In general linear attention mechanisms, a portion of the KV buffer formed by the sum of the two can be expressed as:
$\kappa(K_{1}^{T})V_{1}+\kappa(K_{2}^{T})V_{2}=(1+\lambda)\kappa(K_{2}^{T})V_{2}$ (8)
After introducing $\alpha$ as a modulation coefficient for each term, their sum becomes:
$\alpha_{1}\kappa(K_{1}^{T})V_{1}+\alpha_{2}\kappa(K_{2}^{T})V_{2}$ $=\lambda\alpha_{1}\kappa(K_{2}^{T})V_{2}+\alpha_{2}\kappa(K_{2}^{T})V_{2}$ (9)
$=\alpha_{2}(1+\frac{\lambda\alpha_{1}}{\alpha_{2}})\kappa(K_{2}^{T})V_{2}$
Since in our setup the value of $\alpha_{1}$ and $\alpha_{2}$ are calculated based on the attention scores between global query and keys, it changes with the input samples and the training process. This results in more varied coefficients, meaning that the new matrix composed of these two linearly correlated matrices has a broader range of possible values and greater flexibility. This makes the KV buffer $\sum_{j=1}^{N}\alpha_{j}\kappa(K_{j}^{T})V_{j}$ more likely to achieve a full-rank state.
As previously mentioned (Eq. 8 and Eq. 9), two linearly correlated matrices can be summed to form a single matrix with coefficients. Therefore, we consider here only the case where all matrices are linearly independent. Consider the equation:
$\sum_{j=1}^{N}c_{j}\kappa(K_{j}^{T})V_{j}=C$ (10)
Here, $C\in\mathbb{R}^{d\times d}$ is a full-rank matrix. Since all $\kappa(K_{j}^{T})V_{j}$ are linearly independent, this equation has a unique solution or no solution. After decomposing $c_{j}$ into $d_{j}\alpha_{j}$, the original equation becomes:
$\sum_{j=1}^{N}d_{j}\alpha_{j}\kappa(K_{j}^{T})V_{j}=C$ (11)
Due to the presence of different scalars, the solutions can be multiple, giving the matrix representation a broader scope. This implies that there are more solutions that enable the KV buffer to achieve a full-rank state.
More choices about $\alpha$.
We choose the attention scores between the keys and the global query as the value for $\alpha$, and here we experiment with different selections. When $\alpha$ is set as a learnable vector, its fixed shape of $1\times N$ makes it challenging to apply to higher resolution tasks such as object detection. Therefore, we only evaluate its performance on image classification tasks.
Model Params(M) FLOPs(G) Acc(%)
DeiT-T 6 1.1 72.2
attn score 6 1.1 75.1
learnable 6 1.1 73.8(-1.3)
Table 10: More choice about $\alpha$.
We conduct experiments using the DeiT-T [55] configuration, and the results are shown in Tab. 10. While the learnable $\alpha$ still significantly enhances the model’s performance, its effectiveness is not as impressive as the $\alpha$ based on the attention score.
Appendix B More Analysis About $\phi(.)$
An analysis of the rank-boosting effect.
Consider two matrices $A$ and $B$, $A,B\in\mathbb{R}^{m\times n}$, and represent these matrices as a linear combination of rank-one matrices. That is:
$A=\sum_{i=1}^{r}u_{i}v_{i}^{T},\quad B=\sum_{j=1}^{s}x_{j}y_{j}^{T}$ (12)
where $r={\rm Rank}(A)$, $s={\rm Rank}(B)$, $u_{i},x_{j}\in\mathbb{R}^{m\times 1}$, $v_{i},y_{j}\in\mathbb{R}^{n\times 1}$. Consider the Hardmard product:
$(A\odot B)_{ij}=A_{ij}\odot B_{ij}$ (13)
we can rewrite it as:
$A\odot B$ $=(\sum_{i=1}^{r}u_{i}v_{i}^{T})\odot(\sum_{j=1}^{s}x_{j}y_{j}^{T})$ (14)
$=\sum_{i=1}^{r}\sum_{j=1}^{s}(u_{i}\odot x_{j})(v_{i}\odot y_{j})^{T}$
This expression shows that $A\odot B$ can be viewed as a linear combination of $rs$ rank-one matrices. Therefore, the total rank of $A\odot B$ does not exceed the number of matrices, which is:
${\rm Rank}(A\odot B)\leq rs={\rm Rank}(A)\times{\rm Rank}(B)$ (15)
The above expression effectively demonstrates that when the ranks of the two matrices, $r$ and $s$, are both small, the Hadamard product can effectively raise the upper bound of the matrix rank. Therefore, when augmenting the rank of the output features matrix, the choice of $\phi(.)$ becomes less important; what matters is the use of the Hadamard product. This is consistent with the conclusions we obtained from the ablation experiments in the main text. Regardless of what $\phi(.)$ is, the presence of $\phi(.)$ will always enhance the rank of the matrix.
Figure 9: Overall architecture of RAVLT.
Appendix C Implementation Details
Overall Architectures.
Similar to previous works [19, 35, 18], RAVLT employs a four-stage architectural design, as shown in Fig. 9.
Image Classification.
We adopt the training strategy proposed in DeiT [55] with the only supervision is classification loss. Specifically, all models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. Adam is used with a cosine decay learning rate scheduler and 5 epochs of linear warm-up. The initial learning rate, weight decay, and batch-size are set to 0.001, 0.05, and 1024, respectively. We apply the same data augmentation and regularization as DeiT [55] (RandAugment [8] (randm9-mstd0.5-inc1) , Mixup [72] (prob = 0.8), CutMix [71] (prob = 1.0), Random Erasing (prob = 0.25)).
Object Detection and Instance Segmentation.
We apply RetinaNet [42], Mask-RCNN [29], and Cascaded Mask R-CNN [3] as the frameworks based on the MMDetection [5] to evaluate our models. The models are trained under “1 $\times$” (12 training epochs) and “3 $\times$ +MS” (36 training epochs with multi-scale training) settings. For the “1 $\times$” setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. For the “3 $\times$ +MS”, multi-scale training strategy is applied to randomly resize the shorter side between 480 to 800 pixels. We use the initial learning rate of 1e-4. For RetinaNet, we set the weight decay to 1e-4. For Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.
Semantic Segmentation.
we implement UperNet [64] and SemanticFPN [36] based on MMSegmentation [7] to validate the models. For UperNet, we follow the previous setting [45] and train the model for 160k iterations with the input size of $512\times 512$. For SemanticFPN, we also use the input resolution of $512\times 512$ but train the models for 80k iterations.